a) Verify that B is a basis for IR3 (b) Use the Gram-Schmidt process to produce an orthogonal basis for R (c) Normalize the vectors to produce an orthonormal basis for R3.
Answer #1

2.(a). Let v1 = (1,-1,0)T, v2 = (2,0,1)T and v3 = (5,1,0)T.Let A be the matrix with the given vectors as columns. To determine whether the vectors in B are linearly independent and span R3, we will reduce A toits RREF which is I3. It implies that the given vectors are linearly independent and span R3 so that B is a basis for R3.

(b). Let u1=v1=(1,-1,0)T, u2=v2–proju1(v2)= v2–[(v2.u1)/(u1.u1)]u1 = v2–[(2+0+0)/(1+1+0)]u1 =(2,0,1)T -(1,-1,0)T = (1,1,1)T and u3=v3–proju1(v3)- proju2(v3)= v3–[(v3.u1)/(u1.u1)]u1-[(v3.u2)/(u2.u2)]u2 = v3–[(5-1+0)/(1+1+0)]u1 –[ (5+1+0)/(1+1+1)]u2 =(5,1,0)T -2(1,-1,0)T-2(1,1,1)T =(1,1,-2)T.Then {u1,u2,u3} is an orthogonal basis for R3.

Now, let e1 = u1/||u1||= (1/?2,- 1/?2,0)T, e2 = u2/||u2||= (1/?3,1/?3,1/?3)T and e3 = u3/||u3||= (1/?6,1?6,-2/?6)T. Then {e1,e2,e3} = {(1/?2,- 1/?2,0)T, (1/?3,1/?3,1/?3)T , (1/?6,1?6,-2/?6)T } is an orthonormal basis for R3.